OpenStudy (anonymous):

Triangle ABC has side lengths 7, 24, and 25. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (hoblos):

24^2 + 7^2 =625 25^2= 625=24^2 + 7^2 thus.. yes the side lengths form a Pythagorean triple

OpenStudy (anonymous):

Assume the largest number 25 as length of hypo: then take sqrt ( 7^2 + 24^2) = 25 Since the result matches with the given hypo => Pythagorean triple!